# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ 2.3 Matrix completion with heavy tailed noise
Heavy-tailed data is ubiquitous in many fields such as finance, microeconomics and biology. To address heavy-tailed noise in the context of matrix completion, several estimators have been proposed. Assuming that the second moment of the noise is finite, Minsker [2018] proposes a two-step procedure, where a truncated-type estimator $\tilde{A}$ is obtained first and then a nuclear norm penalized regression is performed on $\tilde{A}$. Fan et al. [2021] also proposes a two-step procedure while a stronger moment condition $E\xi_{i}^{2k}<\infty$ for $k>1$ is required. Elsener and van de Geer [2018] considers a class of loss functions including Huber loss function. They assume the distribution of the noise is symmetric and some additional conditions hold. In particular, Yu et al. [2024] only assumes that the noise variables have finite second moments and still derives strong properties for the estimator under study. The estimator proposed in their paper is based on the Huber loss function, which is one of the fundamental tools for addressing heavy tailed noise. We first define
@@
and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The derivative of $l_{\tau}(\cdot)$ is denoted by
$\psi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
The estimator to be studied is
$\hat{A}_{H}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{1}\right\}.$ (10)
We formally propose the following assumption.
Assumption 3.
$\xi_{i}$ have finite second moments, namely
$E[\xi_{i}^{2}]=\sigma^{2}.$
@@
Let Assumptions 2 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}$ are symmetric. If we take
@@
$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(\sigma^{2},a^{2})rM}{n}.$ (11)
@@
$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
provided that $n\geq C_{4}m\log d$. Here, $C_{1},\ldots,C_{3}$ are universal constants.
We have the following remarks for the comparison of Theorems 5 and 6.
@@
    Theorem 5 improves over Theorem 6 in several aspects: (a) Theorem 5 allows for a more general sampling distribution; (b) It drops the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via the peeling by infinity and nuclear norm to the Huber loss function; (c) Like in Theorems 1 and 3, the $\log d$ factor is removed by utilizing the advanced matrix concentration inequalities in Brailovskaya and van Handel [2024].
@@
    The lower bound in (4) is derived assuming that the noises are Gaussian. Since the Gaussian distribution is of finite second moments, the lower bound also applies to the heavy tailed case. Thus, our new bound (11) together with (4) proves the minimax optimality for the estimator (10).
@@
    Theorem 5 requires symmetry of the noise, while Theorem 6 does not. If the noise is not symmetric and the sampling distribution is uniform, the bound (11) continues to hold under the additional condition $M\geq m\log^{5}d$. While this condition is very mild, this seems to be a pure technical condition and is expected to be unnecessary under future investigation. See our analysis in the proof of (11).
2.4 Matrix completion under corruption
In real applications, the true matrix $A_{0}$ may be corrupted. We consider the situation where the observed noisy entry is from a corrupted matrix $A_{0}=L_{0}+S_{0}$ where $L_{0}$ is of rank $r$ and $S_{0}$ is a sparse corruption matrix. Klopp et al. [2017] considers two corruption models: entrywise corruption and columnwise corruption. We adopt the notation of Klopp et al. [2017]. The total number of observations is denoted by $N$. The set of uncorrupted observations is denoted by $\Omega$, namely $Y_{i}=\langle X_{i},L_{0}\rangle+\xi_{i}$ for $i\in\Omega$. The set of corrupted observations is denoted by $\tilde{\Omega}$, namely $Y_{i}=\langle X_{i},L_{0}+S_{0}\rangle+\xi_{i}$ for $i\in\tilde{\Omega}$. We denote $n=|\Omega|$, $N-n=|\tilde{\Omega}|$, and $\alpha=\frac{N}{n}$.
2.4.1 Entrywise corruption
For entrywise corruption, the set of corrupted entries is denoted by
$\mathcal{I}_{1}=\{(j,k):S_{0}(j,k)\neq 0\},$
where $S_{0}(j,k)$’s are entries of the corruption matrix $S_{0}$. We assume that $S_{0}$ has at most $s$ non-zero entries, i.e. $|\mathcal{I}_{1}|\leq s$. We also need the following assumption.
Assumption 4.
For $i\in\Omega$ and $i\in\tilde{\Omega}$, the sampling is uniform:
$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1},i\in\tilde{\Omega}$
and
$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}m_{2}-|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1}^{c},i\in\Omega.$
Klopp et al. [2017] considers general sampling distributions. To ease the notation and demonstrate more clearly the ideas of our method, we focus on the uniform sampling distribution in Assumption 4. Our results can be easily extended to the general distribution case. For this corruption model, Klopp et al. [2017] considers the following estimator:
$(\hat{L}_{E},\hat{S}_{E})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{1,*}\right\}.$ (12)
We obtain the following theorem.
@@ 2.4.1 Entrywise corruption
Let Assumptions 1 and 4 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$ and $N\leq m_{1}m_{2}$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\Bigg{(}\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\Bigg{)}.$ (13)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] gives the following result.
Theorem 8 (Corollary 11 of Klopp et al. [2017]).
Let Assumptions 1 and 4 hold. Assume that $n\geq C_{1}m\log d\log^{2}m$ and $N\leq m_{1}m_{2}$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\right).$ (14)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] also obtains the minimax lower bound under this setting assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$:
$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{1}m_{2}}\right)\right)\geq C.$ (15)
Here, $C\in(0,1)$, the infimum is taken over all possible estimators for $(L_{0},S_{0})$, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{1}|\leq s$. Our new result removes the $\log d$ factor in the term $\frac{rM\log d}{n}$ in (14). When $rM\geq|\tilde{\Omega}|\log d$, namely the number of corrupted observations is not too large, (13) and (15) together show that the estimator (12) is minimax rate optimal.
2.4.2 Columnwise corruption
For columnwise corruption model, we assume there are at most $s$ corrupted columns, namely $|\mathcal{I}_{2}|\leq s$, where $\mathcal{I}_{2}=\{1\leq i\leq m_{2};S_{0}^{i}\neq\textbf{0}\in\mathbb{R}^{m_{1}}\}$ includes the indices of non-zero columns of $S_{0}$. And we make the following assumption.
Assumption 5.
For $i\in\Omega$, the sampling distribution is uniform:
$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}(m_{2}-|\mathcal{I}_{2}|)}.$
For $i\in\tilde{\Omega}$, the sampling is uniform:
$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}|\mathcal{I}_{2}|}.$
Klopp et al. [2017] proposes the following estimator:
@@ 2.4.2 Columnwise corruption
$(\hat{L}_{C},\hat{S}_{C})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{2,1}\right\}.$ (16)
We obtain the following theorem.
@@
Let Assumptions 1 and 5 hold. Assume $n\geq 2\log d\max(m_{2},m\log^{2}m)$, $N\leq m_{1}m_{2}$, and $n\geq C_{1}m\log^{6}d\log^{2}n$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{Nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
with probability at least $1-\frac{6}{d}$, we have
$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right).$ (17)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] obtains the following result.
Theorem 10 (Corollary 7 in Klopp et al. [2017]).
Let Assumptions 1 and 5 hold. Assume $N\leq m_{1}m_{2}$, and $n\geq 2\log d\max(m_{2},m\log^{2}m)$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{Nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right),$ (18)
where $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
According to Theorem 3 in Klopp et al. [2017], we have the following minimax lower bound for columnwise corruption assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$
$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P_{A_{0}}\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{2}}\right)\right)\geq C,$ (19)
where $C\in(0,1)$ is a universal constant, the infimum is taken over all the estimators, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{2}|\leq s$. We remove the $\log d$ factor attached to $\frac{rM}{n}$ in (18). Comparing (19) and (17), we can conclude that the estimator (16) is minimax rate optimal provided that $rM\geq|\tilde{\Omega}|\log d$, i.e. the set of corrupted observations is not too large.

### v3

@@ 2.1 Matrix completion with heavy tailed noise
We start with our improvement on a recent advance in matrix completion [Yu et al., 2024]. Heavy-tailed data is ubiquitous in many fields such as finance, microeconomics and biology. To address heavy-tailed noise in the context of matrix completion, several estimators have been proposed. Assuming that the second moment of the noise is finite, Minsker [2018] proposes a two-step procedure, where a truncated-type estimator $\tilde{A}$ is obtained first and then a nuclear norm penalized regression is performed on $\tilde{A}$. Fan et al. [2021] also proposes a two-step procedure while a stronger moment condition $E\xi_{i}^{2k}<\infty$ for $k>1$ is required. Elsener and van de Geer [2018] considers a class of loss functions including Huber loss function. They assume the distribution of the noise is symmetric and some additional conditions hold. In particular, Yu et al. [2024] only assumes that the noise variables have finite second moments and still derives strong properties for the estimator under study. The estimator proposed in their paper is based on the Huber loss function, which is one of the fundamental tools for addressing heavy tailed noise.
We first define
@@
and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The estimator to be studied is
$\hat{A}_{H}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{*}\right\}.$ (4)
@@
Let Assumptions 1 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}|X_{i}$ are symmetric. If we take
@@
$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(a^{2},\sigma^{2})rM}{n}.$ (5)
@@
$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
provided that $n\geq C_{4}m\log d$. Here, $C_{1}\sim C_{3}$ are universal constants.
Theorem 1 improves over Theorem 2 in several aspects.
@@
    Our new result eliminates the $\log d$ factor in the convergence rate. It is achieved by utilizing new advanced concentration inequalities in Brailovskaya and van Handel [2024]. See our Lemma 14 for a detailed analysis. Koltchinskii et al. [2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
    $\inf_{\hat{A}}\sup_{A_{0}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (6)
@@
    Theorem 1 allows for a more general sampling distribution while Theorem 2 requires uniform distribution. We also eliminate the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via a different peeling argument [Wainwright, 2019] in the analysis of certain empirical processes.
@@
    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$, which sheds light on the practice of parameter tuning.
  4. 4.
    Theorem 1 requires symmetry of the noise, while Theorem 2 does not. If the noise is not symmetric and the sampling distribution is uniform, the bound (5) continues to hold under the additional condition $M\geq m\log^{5}d$. While this condition is very mild, this seems to be a pure technical condition and is expected to be unnecessary under future investigation. Moreover, assuming that the $\xi_{i}$ have a finite moment of order $2+\kappa$ for some $\kappa>0$, a $\log d$-free bound can be established under a slightly weaker condition on sample size:
    $n\geq\max\left(m\log^{4+\frac{1}{\kappa+1}}d,\log^{\frac{4}{\kappa}}d\right).$
    See our analysis in the proof of inequality (5).
2.2 Matrix completion with known variance
The setting that the noise have distributions with exponential decay is most studied, see, e.g., Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014]. Under the assumption of light tailed noise, an estimator proposed in Klopp [2014] is:
$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$ (7)
We now present our new result:
@@ 2.2 Matrix completion with known variance
Let Assumptions 2 and 3 hold. Assume that $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. If we take
$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
the following bound holds with probability at least $1-\frac{2}{d}$
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(a^{2},\sigma^{2})rM}{n}.$ (8)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 2-3.
To compare with existing bounds, we mention the following theorem:
Theorem 4 (Theorem 7 in Klopp [2014]).
Let Assumption 2, Conditions (1)-(3) in Assumption 3 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
the following bound holds with probability $1-\frac{3}{d}$:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\}.$
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 2-3.
Comparing Theorems 3 and 4, we make several remarks:
  1. 1.
    Theorem 3 eliminates a $\log d$ factor in the convergence rate. Comparing (8) and (6), the removal of the $\log d$ factor establishes the minimax optimality of estimator (7).
  2. 2.
    Theorem 3 also provides the correct order of tuning parameter $\lambda$ for (7). The old one is $O(\sqrt{\log d/(nm)})$ while the correct one in Theorem 3 is $O(\sqrt{1/(nm)})$. The $\log d$ difference can be significant in high dimensional settings.
  3. 3.
    Theorem 4 involves an additional nuisance term $O(\sqrt{\log d/n})$, similarly to Theorem 2. In the high- dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one. However, in less challenging scenarios in which $n$ is sufficiently large, the nuisance term dominates. This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity. In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm. As a result, our new bound does not include the nuisance term $O(\sqrt{\log d/n})$.
  4. 4.
    Our new bound requires a slightly stronger condition on $n$. The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024]. In the most interesting regime where $n\leq m_{1}m_{2}$, the sample size requirement in Theorem 3 becomes $n\geq Cm\log^{5}d$ which is in the typical form of $n\geq m\text{Poly}(\log d)$ in the matrix completion literature. Under both with-replacement and without-replacement settings, the order of $\text{Poly}(\log d)$ often vary with problem assumptions and methods used; see, for example, Koltchinskii et al. [2011], Klopp [2014], Chen et al. [2020]. In our results, the leverage of the concentration inequalities [Brailovskaya and van Handel, 2024] causes a small change of the order of $\text{Poly}(\log d)$. This fluctuation is acceptable in light of the substantial gain in sharpness of the resulting bounds. We emphasize that previous concentration inequalities cannot yield bounds as sharp as ours, even when the sample size is fixed at the same level as in our results.
2.3 Matrix completion with unknown variance of the noise
@@
$\hat{A}_{S}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{*}\right\}.$ (9)
We obtain a new result for (9).
@@
Let Assumptions 2, 3 hold. Let $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. We can take
$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
and with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(a^{2},\sigma^{2})rM}{n},$ (10)
provided that $n\geq 16\mu Mr$. Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 2-3.
We can compare the above bound with the one obtained in Klopp [2014].
Theorem 6 (Theorem 10 in Klopp [2014]).
Let Assumption 2 and and Conditions (1)-(3) in Assumption 3 hold. We can take
$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
and the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
provided that $n\geq 8C_{1}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 2-3.
Comparing the convergence rate in Theorem 6 and the lower bound in (6), they are not matched due to the existence of the $\log d$ factor. Our new bound in (10) resolves the issue and shows that the estimator (9) is indeed minimax rate optimal. Regarding more comparisons between Theorems 5 and 6, we refer to remarks after Theorem 4 for similar comments.